OpenStudy (anonymous):

What is the solution to the equation 21x + 3 = 14 ? Please Explain x ≈ –2.133 x ≈ –1.846 x ≈ 3.867 x ≈ 4.154

OpenStudy (anonymous):

21x + 3 = 14 - 3 -3 21x = 11 /21 /21 x=11/21

OpenStudy (anonymous):

what would that be? sorry, my calculator isn´t working

OpenStudy (anonymous):

wolframalpha.com

OpenStudy (anonymous):

21x+3=14 so you subtract 3 from both sides because you want to get x alone x equals to 11/21

OpenStudy (anonymous):

because 11/21 is a fraction you would keep it that way because it is an exact answer if the want an approx answer you would get 0.5238

OpenStudy (anonymous):

^does that make since?

OpenStudy (anonymous):

i got, 0.523. but its not there

OpenStudy (anonymous):

yes that is also equal to 11/21

OpenStudy (anonymous):

11/21 would be the exact answer 0.523 is the approx answer

OpenStudy (anonymous):

which one would it be closer to? a,b,c or d?

OpenStudy (anonymous):

@jasmiemarie